OpenStudy (anonymous):

A fireworks company has two types of rockets called Zinger 1 and Zinger 2. The polynomial gives the height in feet of Zinger 1 at t seconds after launch. The polynomial: -16t^2 + 165t gives the height of Zinger 2 at t seconds after launch. If the rockets are launched at the same time and both explode 6 seconds after launch, how much higher is Zinger 2 than Zinger 1 when they explode? A. 414 ft B. 990 ft C. 90 ft D. 324 ft

OpenStudy (kropot72):

Do you know how to substitute 6 for t in each equation?

OpenStudy (anonymous):

No, Can U Please Show Me How /: ?

OpenStudy (kropot72):

Substituting the value t = 6 in the first equation gives: (-16 * 36) + (165 * 6) = -576 + 990 = 414 Do you want to try substituting the value t = 6 in the second equation ?

OpenStudy (kropot72):

By the way what is the other polynomial? There is only one in your question.

OpenStudy (kropot72):

I mean the polynomial for Zinger 1

OpenStudy (anonymous):

Well This Is How The Question Was Asked, So I Was Very Confused, But I Think I Got It Now! Thanks For Ur Explanation (:

OpenStudy (kropot72):

You can't get an answer unless you have the other polynomial.

OpenStudy (anonymous):

i dont get it i did the same thing but i got it wrong some how